# A dimensional mass transference principle from balls to open sets and applications to dynamical Diophantine approximation

### v1

@@ Abstract.
The mass transference principle of Beresnevich and Velani is a powerful mechanism for determining the Hausdorff dimension/measure of $\limsup$ sets that arise naturally in Diophantine approximation. However, in the setting of dynamical Diophantine approximation, this principle often fails to apply effectively, as the radii of the balls defining the dynamical $\limsup$ sets generally depend on the orbit of the point $x$ itself.
In this paper, we develop a dimensional mass transference principle that enables us to recover and extend classical results on shrinking target problems, particularly for the $\beta$-transformation and the Gauss map. Moreover, our result shows that the corresponding $\limsup$ sets have large intersection properties. A potentially interesting feature of our method is that, in many cases, shrinking target problems are closely related to finding an appropriate Gibbs measure, which may reveal new aspects of the link between thermodynamic formalism and dynamical Diophantine approximation.
@@ 1. Introduction
The central question in Diophantine approximation is: how well can a given real number $x\in[0,1)$ be approximated by rational numbers. Dating back to Dirichlet, a consequnce of his famous theorem is that for any $x\in[0,1)$,
@@
Throughout, the symbols $\ll$ and $\gg$ will be used to indicate an inequality with an unspecified positive multiplicative constant. If $a\ll b$ and $a\gg b$, we write $a\asymp b$ and say that the quantities $a$ and $b$ are comparable. Let $X$ be a compact metric space equipped with a non-atomic probability measure $\mu$. Suppose there exists a constant $\delta>0$ such that
@@
where the implied constant does not depend on $x$ and $r$. Such a measure is said to be $\delta$-Ahlfors regular.
@@
where $f:[0,1)\to\mathbb{R}$ is a positive function and
@@
is the Birkhoff sum of $f$ along the orbit of $x$. It is well-known that
@@
where $s$ satisfies $P(-s(f+\log 2),T_{2})=0$. Here, $P(\cdot,T_{2})$ denotes the pressure function, see (4.1) for the definition.
A natural question is whether the mass transference principle stated in Theorem 1.1 can be applied to obtain the Hausdorff dimension of $W(T_{2},f,x_{0})$. While the principle is applicable in this setting, it may fail to yield the desired lower bound. This limitation arises because, unlike in classical Diophantine approximation, the targets here are dynamically defined and their radii depend on the orbit of $x$ itself. To apply the principle effectively, one would need to enlarge these balls by a power significantly larger than expected to obtain a $\limsup$ set with full measure. This is precisely why the classical mass transference principle does not directly provide the desired dimension estimates.
@@
To address this shortfall, Wang and Zhang [41] developed an alternative mass transference principle from a dynamical perspective. Utilizing their principle, they successfully recovered the Hausdorff dimension of $W(T_{2},f,x_{0})$. Although the dimension result for $W(T_{2},f,x_{0})$ has been known, their work is significant in providing a new framework that connects shrinking target problems with mass transference principle. However, their principle does not extend to more general transformations such as the $\beta$-transformation or the Gauss map, nor can it be applied to settings where targets are defined by arbitrary open sets rather than balls. The main goal of this paper is to address precisely this issue. Our purpose is to develop a framework capable of handling shrinking target problems — along with various generalizations — for the $\beta$-transformation and the Gauss map, and to extend the theory beyond the classical setting of balls to more general open sets.
@@
(1.3) $\begin{split}&\nu\Big{(}\limsup_{n\to\infty}B(x_{n},r_{n})\Big{)}=1\\ \Longrightarrow\quad&\operatorname{\dim_{H}}\Big{(}\limsup_{n\to\infty}B(x_{n},r_{n}^{\tau})\Big{)}\geq\frac{\operatorname{\dim_{H}}\nu}{\tau}\quad\text{ for $\tau>1$}.\end{split}$
@@
$\underline{D}(\nu,x):=\liminf_{r\to 0}\frac{\log\nu(B(x,r))}{\log r},$
@@
If $\operatorname{\underline{\dim}_{H}}\nu=\operatorname{\overline{\dim}_{H}}\nu$, their common value is denoted by $\operatorname{\dim_{H}}\nu$. However, the quasi-Bernoulli property holds for the Gibbs measures associated with the doubling map, but generally fails for the $\beta$-transformation and the Gauss map. This limitation motivates the development of new concepts capable of handling these cases. To this end, we introduce the notion of quasi-self-conformality of a measure.
@@
    $C^{-1}\dfrac{|x-y|}{|F_{n}|}\leq|f_{n}(x)-f_{n}(y)|\leq C\dfrac{|x-y|}{|F_{n}|}$ for all $x,y\in F_{n}$, where $|A|$ denotes the diameter of a set $A$;
@@
    $C^{-1}\nu(A)\leq\nu^{(n)}(A)\leq C\nu(A)\quad\text{for any Borel set $A$}.$
@@
Remark 1.
Definition 1.2 (1) implies that:
  1. (a)
    For any $x\in F_{n}$ and $0<r<|F_{n}|$,
    $B(f_{n}(x),C^{-1}r/|F_{n}|)\subset f_{n}(B(x,r)\cap F_{n})\subset B(f_{n}(x),Cr/|F_{n}|).$
  2. (b)
    If $X$ supports a $\delta$-Ahlfors regular measure $\mu$, then there exists an absolute constant $c\geq 1$ such that for any $F_{n}$,
    $F_{n}\subset B(x_{n},c|F_{n}|)\mathrm{\quad and\quad}c^{-1}|F_{n}|^{\delta}\leq\mu(F_{n})\leq c|F_{n}|^{\delta},$
    where $x_{n}\in F_{n}$.
To formulate our main result, we recall the notion of Hausdorff content. In this paper, we focus on the case where the ambient space $X$ is a compact subset of $\mathbb{R}^{d}$. For any $s\geq 0$ and a set $A$, the $s$-dimensional Hausdorff content of $A$ is defined by
$\mathcal{H}_{\infty}^{s}(A)=\inf\bigg{\{}\sum_{i}|B_{i}|^{s}:A\subset\bigcup_{i\geq 1}B_{i},\text{ where $B_{i}$ are balls}\bigg{\}}.$
@@
Let $0<s\leq\operatorname{\dim_{H}}X$. We define $\mathscr{G}^{s}(X)$ to be the class of $G_{\delta}$-subsets $A$ of $X$ such that there exists a constant $c>0$ such that for any $0<t<s$ and any ball $B$,
(1.4) $\mathcal{H}_{\infty}^{t}(A\cap B)>c\mathcal{H}_{\infty}^{t}(B).$
If $X$ supports a $\delta$-Ahlfors regular measure, then the class $\mathscr{G}^{s}(X)$ is closed under countable intersections, and moreover,
$\operatorname{\dim_{H}}A\geq s\quad\text{for all $A\in\mathscr{G}^{s}(X)$}.$
@@
Let $X\subset\mathbb{R}^{d}$ be a compact subset equipped with a $\delta$-Ahlfors regular measure $\mu$. Let $\nu$ be a quasi-self-conformal measure with respect to a collection of closed sets $\mathcal{F}=\{F_{n}\}$ in $X$, such that
(1.5) $\mu\Big{(}\limsup_{n\to\infty}F_{n}\Big{)}=1.$
Suppose that there exist a sequence of balls $\{B(x_{n},r_{n})\}$ and a sequence of open sets $\{E_{n}\}$ satisfying the following conditions:
@@
    $\mathcal{H}_{\infty}^{s}(E_{n})\gg r_{n}^{\operatorname{\overline{\dim}_{H}}\mu},$
@@
The assumption that $X\subset\mathbb{R}^{d}$ is essential, as our arguments rely on the Besicovitch covering theorem, which generally does not hold in metric spaces. This arises because the measure $\nu$ is generally not doubling, which prevents many standard covering lemmas from applying effectively — except for the Besicovitch covering theorem. The condition in (1.5) serves to characterize the extent to which the measure $\nu$ exhibits quasi-self-conformality. In many case, a measure $\nu$ satisfy this condition is a Gibbs measure, and thus generally singular to the ambient measure $\mu$.
There are two different but closely related ways to understand the connection between our result and the mass transference principle of Beresnevich and Velani. To set the stage, suppose there exist two sequences of balls $\{B(x_{n},r_{n})\}$ and $\{B(y_{n},t_{n})\}$ such that
$\mu\Big{(}\limsup_{n\to\infty}B(x_{n},r_{n})\Big{)}=1\mathrm{\quad and\quad}\nu\Big{(}\limsup_{n\to\infty}B(y_{n},t_{n})\Big{)}=1,$
where $\mu$ and $\nu$ are measures as in Theorem 1.4. The first perspective may be regarded as a counterpart to (1.3). Fix $\tau>1$, and apply Theorem 1.1 and Theorem 1.4 with $s=\operatorname{\overline{\dim}_{H}}\nu/\tau$, respectively. Note that $\delta=\operatorname{\dim_{H}}\mu$. Then, we obtain
$\operatorname{\dim_{H}}\Big{(}\limsup_{n\to\infty}B(x_{n},r_{n}^{\tau})\Big{)}\geq\operatorname{\dim_{H}}\mu/\tau\mathrm{\quad and\quad}\operatorname{\dim_{H}}\Big{(}\limsup_{n\to\infty}B(y_{n},t_{n}^{\tau})\Big{)}\geq\operatorname{\overline{\dim}_{H}}\nu/\tau,$
and these lower bounds coincide when $\mu=\nu$. In other words, if enlarging the radii of the balls by a power of $1/\tau$ yields a $\limsup$ set of full measure (with respect to $\mu$ or $\nu$), then the $\limsup$ set of the original balls has Hausdorff dimension at least
$\frac{\text{Hausdorff dimension of the measure}}{\tau}.$
The second perspective is expressed as follows:
$\begin{split}&\mathcal{H}_{\infty}^{\delta/\tau}(B(x_{n},r_{n}^{\tau}))\asymp r_{n}^{\delta}\asymp\mu(B(x_{n},r_{n}))\quad\text{and}\quad\mu\Big{(}\limsup_{n\to\infty}B(x_{n},r_{n})\Big{)}=1\\ \Longrightarrow\quad&\operatorname{\dim_{H}}\Big{(}\limsup_{n\to\infty}B(x_{n},r_{n}^{\tau})\Big{)}\geq\delta/\tau,\end{split}$
while
$\begin{split}&\mathcal{H}_{\infty}^{\operatorname{\overline{\dim}_{H}}\nu/\tau}(B(y_{n},t_{n}^{\tau}))\asymp r_{n}^{\operatorname{\overline{\dim}_{H}}\nu}\sim\nu(B(y_{n},t_{n}))\quad\text{and}\quad\nu\Big{(}\limsup_{n\to\infty}B(y_{n},t_{n})\Big{)}=1\\ \Longrightarrow\quad&\operatorname{\dim_{H}}\Big{(}\limsup_{n\to\infty}B(y_{n},t_{n}^{\tau})\Big{)}\geq\operatorname{\overline{\dim}_{H}}\nu/\tau.\end{split}$
Here we use $\sim$’ instead of $\asymp$’ because two quantities in general no comparable. That is, if a $\limsup$ set has full measure (with respect to $\mu$ or $\nu$) and we shrink each ball to a smaller one whose $s$-dimensional Hausdorff content is comparable to the measure of the original, then the resulting $\limsup$ set has Hausdorff dimension at least $s$. Although the two perspectives are equivalent, the latter is more flexible and naturally leads us to consider extending from balls to general open sets.
We now apply Theorem 1.4 to recover and extend classical results for the $\beta$-transformation and the Gauss map. For a Lipschitz function $h$, its Lipschitz constant is defined as the smallest $L>0$ such that for any $x,y\in X$,
@@
Let $\{h_{n}\}$ be a sequence of Lipschitz functions with uniformly bounded Lipschitz constants and let $f$ be a positive function defined on $X$. The modified shrinking target sets are defined as
@@
Suppose that $T$ is either the $\beta$-transformation or the Gauss map. Then,
@@
The lower bound for the Hausdorff dimension of $W(T,f,\{h_{n}\})$ implied in Theorem 1.5 was previously established in [10, 37, 32, 42]. However, those results rely on the construction of large Cantor-type sets and do not imply the large intersection property. Interestingly, Theorem 1.4 offers a different perspective: the problem is reduced to seeking a suitable Gibbs measure and estimating its Hausdorff dimension. This perspective may provide new insights into the interplay between thermodynamic formalism and dynamical Diophantine approximation.
@@
After completing the proofs of our main results, the author became aware that Daviaud [14] had employed some similar ideas to study the shrinking target problem for self-conformal sets with overlaps. However, his results neither imply the large intersection property nor can they be directly applied to the $\beta$-transformation or the Gauss map.
@@
The structure of the paper is as follows. In Section 2, we collect several foundational results and technical tools that will be used throughout the paper. Section 3 is devoted to the proof of our main result, i.e. Theorem 1.4. In Section 4, we review the definition and key properties of the $\beta$-transformation, and apply our main result to obtain the large intersection property of $W(T_{\beta},f,\{h_{n}\})$ in this setting. Section 5 serves a similar purpose for the Gauss map: we recall its basic properties and then apply our theorem to derive the large intersection properties of $W(G,f,\{h_{n}\})$ and $F_{m}(B)$.

### v2

@@ Abstract.
The mass transference principle of Beresnevich and Velani is a powerful mechanism for determining the Hausdorff dimension/measure of $\limsup$ sets that arise naturally in classical Diophantine approximation. However, in the setting of dynamical Diophantine approximation, this principle often fails to apply effectively, as the radii of the balls defining the dynamical $\limsup$ sets generally depend on the orbit of the point $x$ itself.
In this paper, we develop a dimensional mass transference principle that enables us to recover and extend classical results on shrinking target problems, particularly for the $\beta$-transformations and the Gauss map. Moreover, our result shows that the corresponding $\limsup$ sets have large intersection properties. A potentially interesting feature of our method is that, in many cases, shrinking target problems are closely related to finding an appropriate Gibbs measure, which may reveal new aspects of the link between thermodynamic formalism and dynamical Diophantine approximation.
@@ 1. Introduction
The central question in Diophantine approximation is: how well can a given real number $x\in[0,1)$ be approximated by rational numbers. Dating back to Dirichlet, a consequence of his famous theorem is that for any $x\in[0,1)$,
@@
Throughout, the symbols $\ll$ and $\gg$ will be used to indicate an inequality with an unspecified positive multiplicative constant. If $a\ll b$ and $a\gg b$, we write $a\asymp b$ and say that the quantities $a$ and $b$ are comparable. Let $|E|$ denote the diameter of a set $E$. Let $X$ be a compact metric space equipped with a non-atomic probability measure $\mu$. Suppose there exists a constant $\delta>0$ such that for any $x\in X$ and $0<r<|X|$
@@
where the implied constant does not depend on $x$ and $r$. Such a measure is said to be $\delta$-Ahlfors regular. If $X$ supports a $\delta$-Ahlfors regular measure $\mu$, then we have that $\operatorname{\dim_{H}}X=\delta$ and $\mu$ is comparable to the $\delta$-dimensional Hausdorff measure (denoted by $\mathcal{H}^{\delta}$) restricted to $X$ — see [17] for the details.
@@
where $x_{0}\in[0,1]$, $f:[0,1]\to\mathbb{R}_{+}$ is a Hölder continuous function and
@@
is the Birkhoff sum of $f$ along the orbit of $x$. It is already known (see e.g. [43]) that
@@
with $s$ satisfying
(1.3) $P(-s(f+\log 2),T_{2}):=\lim_{n\to\infty}\frac{1}{n}\log\sum_{i=0}^{2^{n}-1}\big(2^{-n}e^{-S_{n}f(y_{n,i})}\big)^{s}=0,$
where $y_{n,i}\in[i2^{-n},(i+1)2^{-n}]$ is chosen arbitrarily, and $P(\cdot,T_{2})$ denotes the pressure function with respect to $T_{2}$ (see Section 4.2 for further details).
A natural question is whether the mass transference principle stated in Theorem 1.1 can be applied to obtain the Hausdorff dimension of $W(T_{2},f,x_{0})$. To explore this, note that $f$ is Hölder continuous, and so $W(T_{2},f,x_{0})$ can be roughly expressed as
$\bigcap_{N=1}^{\infty}\bigcup_{n=N}^{\infty}\bigcup_{i=0}^{2^{n}-1}B(x_{n,i},2^{-n}e^{-S_{n}f(x_{n,i})})=:\bigcap_{N=1}^{\infty}\bigcup_{n=N}^{\infty}E_{n},$
where $x_{n,i}\in[i2^{-n},(i+1)2^{-n}]$ satisfies $T_{2}^{n}x_{n,i}=x_{0}$. If $f$ is not constant, then for a fixed $n\in\mathbb{N}$, the terms $e^{-S_{n}f(x_{n,i})}$ are not identical as $i$ varies. Therefore, $E_{n}$ consists of intervals of varying lengths, some of which may be very small, while others may be quite large (see Figure 2 for an illustration and Figure 1 for a comparison). Although the initial intervals are well-separated, enlarging them alters the situation, which constitutes the most significant difference compared to $W(\tau)$.
Figure 1. When $f$ is constant, $E_{2}$ consists of intervals of equal length.
Figure 2. When $f$ is not constant, $E_{2}$ consists of intervals of varying lengths.
To be able to apply Theorem 1.1, it is necessary to enlarge the intervals in $E_{n}$’s to obtain a larger $\limsup$ set with full one-dimensional Lebesgue measure (denoted by $\mathcal{L}$) restricted to $[0,1]$. It follows from (1.3) that for any $\epsilon>0$ and any sufficiently large $n$ (depending on $\epsilon$),
$\sum_{i=0}^{2^{n}-1}\left(2^{-n}e^{-S_{n}f(x_{n,i})}\right)^{s}=O\left(e^{n\epsilon}\right),$
which means that the total length of the intervals $B\big(x_{n,i},\big(2^{-n}e^{-S_{n}f(x_{n,i})}\big)^{s}\big)$ in
$E_{n,s}:=\bigcup_{i=0}^{2^{n}-1}B\big(x_{n,i},\big(2^{-n}e^{-S_{n}f(x_{n,i})}\big)^{s}\big)$
is quite large. However, this neither guarantees that
(1.4) $0,1]=E_{n,s}\quad\text{nor that}\quad\mathcal{L}\Big(\limsup_{n\to\infty}E_{n,s}\Big)=1,$
because the lengths of these intervals are different, making it possible for them to overlap (see Figure 4 for an illustration and Figure 3 for a comparison).
Figure 3. When $f$ is constant, the enlarged intervals contained in $E_{2,s}$ are well-separated and therefore efficiently cover $[0,1]$.
Figure 4. When $f$ is not constant, the enlarged intervals contained in $E_{2,s}$ may overlap and therefore fail to cover $[0,1]$ efficiently.
Therefore, one should choose a parameter $t$ significantly smaller than $s$ to ensure that one of the equalities in (1.4) holds with $s$ replaced by $t$, i.e., $E_{n,t}$ effectively covers $[0,1]$. This leads to the situation that although Theorem 1.1 is still applicable, it does not directly provide the desired dimension estimates.
@@
To address this shortfall, Wang and Zhang [46] developed an alternative mass transference principle from a dynamical perspective. Utilizing their principle, they successfully recovered the Hausdorff dimension of $W(T_{2},f,x_{0})$. Although the dimension result for $W(T_{2},f,x_{0})$ has been known, their work is significant in providing a new framework that connects shrinking target problems with mass transference principle. However, their principle does not extend to more general transformations such as the $\beta$-transformations or the Gauss map, nor can it be applied to settings where targets are defined by arbitrary open sets rather than balls. The main goal of this paper is to address precisely this issue. Our purpose is to develop a framework capable of handling shrinking target problems — along with various generalizations — for the $\beta$-transformations and the Gauss map, and to extend the theory beyond the classical setting of balls to more general open sets. It is also worth noting that the setting in [46] is somewhat technical — for example, it relies on the dynamical ubiquity and topological exactness assumptions — and is almost entirely different from the current one that we will describe.
@@
(1.5) $\begin{split}&\nu\Big(\limsup_{n\to\infty}B(x_{n},r_{n})\Big)=1\quad\text{with\quad$\lim_{n\to\infty}r_{n}=0$}\\ \Longrightarrow\quad&\operatorname{\dim_{H}}\Big(\limsup_{n\to\infty}B(x_{n},r_{n}^{\tau})\Big)\geq\frac{\operatorname{\dim_{H}}\nu}{\tau}\quad\text{ for any $\tau>1$}.\end{split}$
@@
(1.6) $\underline{D}(\nu,x):=\liminf_{r\to 0}\frac{\log\nu(B(x,r))}{\log r},$
@@
If $\operatorname{\underline{\dim}_{H}}\nu=\operatorname{\overline{\dim}_{H}}\nu$, their common value is denoted by $\operatorname{\dim_{H}}\nu$. However, the quasi-Bernoulli property holds for the Gibbs measures associated with the doubling map, but generally fails for the $\beta$-transformations and the Gauss map. This limitation motivates the development of new concepts capable of handling these cases. To this end, we introduce the notion of quasi-self-conformality of a measure.
@@
    $C^{-1}\dfrac{|x-y|}{|F_{n}|}\leq|f_{n}(x)-f_{n}(y)|\leq C\dfrac{|x-y|}{|F_{n}|}$ for all $x,y\in F_{n}$;
@@
    $C^{-1}\nu(E)\leq\nu^{(n)}(E)\leq C\nu(E)\quad\text{for any Borel set $E$}.$
@@
To formulate our main result, we recall the notion of Hausdorff content. In this paper, we focus on the case where the ambient space $X$ is a compact subset of $\mathbb{R}^{d}$. For any $s\geq 0$ and a set $E\subset\mathbb{R}^{d}$, the $s$-dimensional Hausdorff content of $E$ is defined by
$\mathcal{H}_{\infty}^{s}(E)=\inf\bigg\{\sum_{i}|B_{i}|^{s}:E\subset\bigcup_{i\geq 1}B_{i},\text{ where $B_{i}$ are balls}\bigg\}.$
@@
Let $0<s\leq\operatorname{\dim_{H}}X$. We define $\mathscr{G}^{s}(X)$ to be the class of $G_{\delta}$-subsets $E$ of $X$ such that there exists a constant $c>0$ such that for any $0<t<s$ and any ball $B$,
(1.7) $\mathcal{H}_{\infty}^{t}(E\cap B)>c\mathcal{H}_{\infty}^{t}(B).$
Falconer originally defined this property in $\mathbb{R}^{d}$, and it has later been extended to general metric spaces. It has been shown, for example in [22, Theorem 2.4], that if $X$ supports a $\delta$-Ahlfors regular measure, then the class $\mathscr{G}^{s}(X)$ is closed under countable intersections, and moreover,
$\operatorname{\dim_{H}}E\geq s\quad\text{for all $E\in\mathscr{G}^{s}(X)$}.$
@@
Let $X\subset\mathbb{R}^{d}$ be a compact subset equipped with a $\delta$-Ahlfors regular probability measure $\mu$. Let $\nu$ be a quasi-self-conformal probability measure with respect to a collection of closed sets $\mathcal{F}=\{F_{n}\}$ in $X$, such that
(1.8) $\mu\Big(\limsup_{n\to\infty}F_{n}\Big)=1\mathrm{\quad and\quad}\lim_{n\to\infty}|F_{n}|=0.$
Suppose that there exist a sequence of balls $\{B(x_{n},r_{n})\}$ with centers in $X$, and a sequence of open sets $\{E_{n}\}$ in $X$, satisfying the following conditions:
@@
    $\mathcal{H}_{\infty}^{s}(E_{n})\gg r_{n}^{\operatorname{\overline{\dim}_{H}}\nu},$
@@
The assumption that $X\subset\mathbb{R}^{d}$ is essential, as our arguments rely on the Besicovitch covering theorem, which generally does not hold in metric spaces. This arises because the measure $\nu$ is generally not doubling, which prevents many standard covering lemmas from applying effectively — except for the Besicovitch covering theorem.
Remark 2.
The condition in (1.8) serves to characterize the extent to which the measure $\nu$ exhibits quasi-self-conformality. In many case, a measure $\nu$ satisfy this condition is a Gibbs measure, and thus generally singular to the ambient measure $\mu$. For example, if $\mu$ is the Lebesgue measure, and $\nu$ is the (p,1p)(p,1-p)-Bernoulli measure with $0<p<1$, then we can take $F_{n}$ to be the dyadic intervals and $f_{n}$ be the linear map that sends $F_{n}$ onto $[0,1]$. Therefore, $\nu$ satisfies (1.8). We will discuss additional examples of interest, such as the $\beta$-transformations in Section 4 and the Gauss map in Section 5.
Remark 3.
It is important to highlight the differences and the novelty of Theorem 1.4 in comparison with existing ‘balls to open sets’ mass transference principles [22, 31, 38, 48]. First, the results of the latter focus on the cases $\nu=\mu$, in which the estimate $\mu(B(x,r))\asymp r^{\delta}$ for arbitrary balls plays a crucial role in the proof. However, if $\nu$ is singular with respect to $\mu$, then $\nu$ does not satisfy this estimate and, worse still, may fail to be doubling. Therefore, new approaches are required, and this is one of the reasons why we introduce the notion of quasi-self-conformality; for further details, see Remark 8. Second, as illustrated by the example of $W(T_{2},f,x_{0})$, the results in [22, 31, 38, 48] cannot be applied in a straightforward manner to determine the Hausdorff dimension of $W(T_{2},f,x_{0})$. As we will see in Section 4.3, together with some known results from thermodynamic formalism, Theorem 1.4 can be directly applied to obtain the lower bound for $\operatorname{\dim_{H}}W(T_{2},f,x_{0})$, in the same manner that Theorem 1.1 is used to obtain $\operatorname{\dim_{H}}W(\tau)$. This largely demonstrates the effectiveness of our theorem in addressing Diophantine approximation problems in dynamical systems.
We now apply Theorem 1.4 to recover and extend classical results for the $\beta$-transformations and the Gauss map. For a Lipschitz function $h:X\to X$, its Lipschitz constant is defined as the smallest $L>0$ such that for any $x,y\in X$,
@@
Let $\{h_{n}\}$ be a sequence of Lipschitz functions with uniformly bounded Lipschitz constants and let $f$ be a positive and continuous function defined on $X$. The modified shrinking target sets are defined as
@@
Let $f$ be a positive and continuous function defined on $X$. Suppose that $X=[0,1]$ and that $T$ is either the $\beta$-transformation or the Gauss map. Then,
@@
The lower bound for the Hausdorff dimension of $W(T,f,\{h_{n}\})$ implied in Theorem 1.5 was previously established in [10, 41, 34, 47]. However, those results rely on the construction of large Cantor-type sets and do not imply the large intersection property. Interestingly, Theorem 1.4 offers a different perspective: the problem is reduced to seeking a suitable Gibbs measure and estimating its Hausdorff dimension. This perspective may provide new insights into the interplay between thermodynamic formalism and dynamical Diophantine approximation. It can be seen from our proof that Theorem 1.4 is also applicable to the shrinking target problems in expanding Markov maps with finite partitions on $[0,1]$, see Remark 10 for further details. The only reason we restrict ourselves to these two cases is that the $\beta$-transformations and the Gauss map are among the most well-known non-Markov map and expanding map with infinitely many branches, respectively.
@@
After completing the proofs of our main results, the author became aware that Daviaud [14] had employed some similar ideas to study the shrinking target problem for self-conformal sets with overlaps. However, his results neither imply the large intersection property nor can they be directly applied to the $\beta$-transformations or the Gauss map.
@@
The structure of the paper is as follows. In Section 2, we collect several foundational results and technical tools that will be used throughout the paper. Section 3 is devoted to the proof of our main result, i.e. Theorem 1.4. In Section 4, we review the definition and key properties of the $\beta$-transformations, and apply our main result to obtain the large intersection property of $W(T_{\beta},f,\{h_{n}\})$ in this setting. Section 5 serves a similar purpose for the Gauss map: we recall its basic properties and then apply our theorem to derive the large intersection properties of $W(G,f,\{h_{n}\})$ and $F_{m}(B)$.